OpenStudy (anonymous):

QUICK HELP! When "subtracting" two negative numbers, how do you know which one to turn into a positive number hence that you can subtract? what if you had to subtract both numbers and order didnt matter such as in the algebra problem -2+x=-3+x. How do you subtract, or add like terms in such manner?

OpenStudy (anonymous):

If you want to solve for x, that would be: -2 + x = -3 + x -2 + 2 + x = -3 + x + 2 0 + x = -1 + x x - x = -1 0 does not equal -1, therefore there exists no solution for this.

OpenStudy (anonymous):

Now to answer your question, suppose we have: 

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

You turn the SECOND number into positive when subtracting, so what happens we subtract is that you can think of this sign change occurring: